OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point m=9,(3,2)

OpenStudy (anonymous):

use point slope form \[\Large y-y_{0}=m(x-x_{0})\] \[\Large y-2=9(x-3)\]

OpenStudy (anonymous):

ya i have not used that.

OpenStudy (anonymous):

so its ok now ?? or do you need further help ?